# Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks

### v1

@@ Abstract
We show that even the most recent safety-aligned LLMs are not robust to simple adaptive jailbreaking attacks. First, we demonstrate how to successfully leverage access to logprobs for jailbreaking: we initially design an adversarial prompt template (sometimes adapted to the target LLM), and then we apply random search on a suffix to maximize the target logprob (e.g., of the token “Sure”), potentially with multiple restarts. In this way, we achieve nearly 100% attack success rate—according to GPT-4 as a judge—on GPT-3.5/4, Llama-2-Chat-7B/13B/70B, Gemma-7B, and R2D2 from HarmBench that was adversarially trained against the GCG attack. We also show how to jailbreak all Claude models—that do not expose logprobs—via either a transfer or prefilling attack with 100% success rate. In addition, we show how to use random search on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—which is the algorithm that brought us the first place in the SaTML’24 Trojan Detection Competition. The common theme behind these attacks is that adaptivity is crucial: different models are vulnerable to different prompting templates (e.g., R2D2 is very sensitive to in-context learning prompts), some models have unique vulnerabilities based on their APIs (e.g., prefilling for Claude), and in some settings it is crucial to restrict the token search space based on prior knowledge (e.g., for trojan detection). We provide the code, prompts, and logs of the attacks at https://github.com/tml-epfl/llm-adaptive-attacks.
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) carry the inherent risk of misuse, such as producing toxic content, spreading misinformation or supporting harmful behaviors. To mitigate these risks, safety alignment is commonly employed—a fine-tuning phase where models are guided to generate responses judged safe by humans and to refuse responses to potentially harmful queries (Bai et al., 2022; Touvron et al., 2023). Although safety alignment is effective in general, several works have shown that it can be circumvented using adversarial prompts. These are inputs specifically designed to induce harmful responses from the model, a practice known as jailbreaking attacks (Mowshowitz, 2022; Zou et al., 2023; Chao et al., 2023).
@@
Jailbreaking attacks vary in their knowledge of the target LLM (ranging from white- to black-box approaches, or API-only access), complexity (involving manual prompting, standard optimization techniques, or auxiliary LLMs), and computational cost. Moreover, the nature of the jailbreaks they produce differs: some methods insert strings with no semantic meaning (Zou et al., 2023), while others rephrase user requests to maintain natural language (Mehrotra et al., 2023). The effectiveness of these attacks can significantly vary, achieving a high success rate on some target models but also drastically failing on others. Finally, some LLMs, such as the Llama-2-Chat family (Touvron et al., 2023), seem to maintain their robustness against these attacks. At the same time, new defensive mechanisms designed to counteract jailbreaks are emerging (Robey et al., 2023; Mazeika et al., 2024).
@@
In this work, we examine the safety of leading safety-aligned LLMs in terms of robustness to jailbreaks. We show that it is feasible to leverage the information available about each model, derived from training details or inference (e.g., logprobs), to construct simple adaptive attacks. Our main tool consists of manually designing a universal template (i.e., a single template is used for all unsafe requests) for each target model (or family of models), supported by random search (RS) (Rastrigin, 1963) when the logprobs of the generated tokens are (at least partially) accessible. Unlike some prior works (Zou et al., 2023; Geisler et al., 2024), our approach does not require gradient information (even for open-weight models) or auxiliary LLMs (Chao et al., 2023; Mehrotra et al., 2023; Zeng et al., 2024) to iteratively optimize the jailbreaks. In this way, using the dataset of unsafe prompts from Chao et al. (2023), we obtain a close to 100% attack success rate on all leading safety-aligned LLMs, including GPT-3.5, GPT-4, Claude-3, Gemma, Llama-2-Chat, and the adversarially trained R2D2, outperforming the existing techniques. We provide a summary of these results in Table 1, and we show an illustrative example of a successful transfer attack on Claude 3 Sonnet in Figure 1. Additionally, we show how to combine manual adaptation and RS for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—enabling us to secure the first place in the SaTML’24 Trojan Detection Competition (Rando & Tramèr, 2024).
@@
Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it becomes evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020); Croce et al. (2022b), distilling key observations from our experiments.
@@ 3.1 Setting
We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 9 for the prompt) in line with the criteria established by Chao et al. (2023), where jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. Additionally, we include results using the rule-based judge from Zou et al. (2023) in the appendix for comparison.
@@ 3.2 Methodology
Our adaptive attacks leverage several key components, which we combine or slightly adapt depending on the target LLMs.
@@
The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We optimized the set of rules one by one on the GPT-3.5 Turbo model to maximize the attack success rate and avoid the built-in safety guardrails. We provide the complete template in Table 2. As we will see in Section 4 (Table 5), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We also designed another prompt template (referred to as the in-context prompt) that includes an example of unsafe behavior which the model is encouraged to imitate (see Table 8).
@@
We use a simple random search (RS) algorithm (Rastrigin, 1963) adapted for jailbreaking language models. The algorithm is as follows:
@@
We use adversarial suffixes that are initialized with $25$ tokens, we use up to 10 0001000010\,00010 000 iterations and potentially a few random restarts. This strategy, including the preference for suffixes over prefixes and the focus on maximizing the log-probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We opted for random search (a) due to its simplicity and efficiency, requiring only scores (such as logprobs) instead of gradients (thus reducing the memory demands), and (b) motivated by its success in adversarial attacks on vision models (Andriushchenko et al., 2020; Croce et al., 2022a). We provide further details in the code. [^1] https://github.com/tml-epfl/llm-adaptive-attacks
@@
It is well-known that initialization plays a key role in optimization algorithms, including in RS-based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request as the initialization for RS on more challenging requests. We refer to this approach as self-transfer. It significantly boosts the attack success rate on some LLMs like Llama-2-Chat, Gemma, and GPT-4.
@@
Some APIs like Claude allow users to directly prefill the LLM’s response with a specified beginning, making the aforementioned optimization procedure unnecessary. In that case, we explore prefilling the response with a string that corresponds to the target behavior (e.g., “Sure, here is how to make a bomb”).
@@ 4 Jailbreaking Leading Safety-Aligned LLMs
In this section, we detail the adaptive attacks we have developed for several families of leading safety-aligned LLMs. We provide a summary of main evaluations here and show the rest in Table 19 in the appendix where we also present results on Vicuna-13B and Mistral-7B.
4.1 Jailbreaking Llama-2 and Gemma models
Here, we focus on open-weights Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023) and Gemma-7B models (Google, 2023). These models have been significantly safety-aligned, rendering them resilient to jailbreaks even in white-box scenarios (Zou et al., 2023).
@@ 4.1 Jailbreaking Llama-2 and Gemma models
The key element to jailbreak Llama-2-Chat models is self-transfer, where successful adversarial suffixes found by RS on simpler requests are used as initialization for RS on more complex requests. Notably, these adversarial strings tend to be to some extent transferable across different model sizes (e.g., from 7B to 13B models), but for the best result we repeat the self-transfer procedure for each model size separately. The same approach is also successful on Gemma-7B, although prompt + RS alone already demonstrates high attack success rate.
@@ Results.
For Llama-2-Chat models, Table 3 shows that our standard adversarial prompt templates yield a 0% success rate, confirming the effectiveness of their safety alignment. When we apply Prompt + RS the attack success rate (ASR) increases to 48%. Ultimately, our composite attack strategy—which combines prompting, random search, and self-transfer—achieves a 100% attack success rate for all LLMs, surpassing all existing methods. For Llama-2-Chat-7B, the best reported success rate is 92% by PAP (Zeng et al., 2024), an LLM-assisted method. However, this method requires 10 restarts to reach such efficacy, and its success rate drops to 46% with only one restarts as in our approach. Meanwhile, for the 13B and 70B models, Mazeika et al. (2024) reports ASR below 40%, while there is no prior evaluation available for Gemma-7B.
Table 3: Llama-2 and Gemma models. We report the attack success rate according to the GPT-4 judge.
@@ 4.2 Jailbreaking R2D2 model
R2D2 uses adversarial training (Madry et al., 2018), a technique effective for obtaining vision models robust to $\ell_{p}$-bounded adversarial perturbations (Madry et al., 2018; Croce et al., 2021), to make LLMs robust to jailbreaks.
@@
Similarly to Llama-2-Chat, the standard prompt template, alone or with RS, shows limited effectiveness. However, in contrast with Llama-2-Chat, self-transfer is ineffective here. Thus, we circumvent safety guardrails using an in-context prompt (see Table 8), which we found the model to be particularly sensitive to. We use random search on top of the in-context prompt to maximize the probability of the initial token “Step” (instead of “Sure”) to be consistent with the new prompt template.
@@ Results.
The limitation of the API providing only the top-5 log-probabilities is not critical, as it is often straightforward to prompt a desired token, like “Sure,” to appear in the top-5. A more challenging issue is the non-deterministic output, since RS does not necessarily have a correct signal to refine the adversarial string. As illustrated in Figure 2, identical queries can yield varying log-probabilities, even with a fixed seed parameter and temperature zero in the API. The randomness makes random search less effective, although it still succeeds to a large extent.
@@ 4.4 Jailbreaking Claude models
Claude models are known for their high safety levels. In line with this, Anthropic does not provide access to logprobs for these models which prevents direct iterative attack like random search. Thus, we first test a transfer attack using an adversarial suffix from GPT-4. We enhance the attack with multiple random restarts to leverage different generations with temperature one. Subsequently, we investigate an attack method utilizing Anthropic’s prefilling feature, [^2] https://docs.anthropic.com/claude/docs/prefill-claudes-response a functionality not commonly available from other LLM providers like OpenAI.
@@
The prefilling feature makes jailbreaking straightforward on Claude models, even without any search (Table 6). For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The latest Claude 2.1 model, released in November 2023, is significantly more robust to both transfer and prefilling attacks. Nonetheless, we are able to get $100\%$ ASR with 100 restarts. We note that GPT-4 as a semantic judge sometimes has false positives, more often so on Claude 2.1. At the same time, the technical report of Claude-3 (Anthropic, 2023) mentions fewer refusals in its release announcement, which, in our findings, correlated with vulnerability to jailbreaking. We provide more complete experimental results, including the number of restarts in each case, in Tables 16 and 17 in the appendix.

### v2

@@ Abstract
We show that even the most recent safety-aligned LLMs are not robust to simple adaptive jailbreaking attacks. First, we demonstrate how to successfully leverage access to logprobs for jailbreaking: we initially design an adversarial prompt template (sometimes adapted to the target LLM), and then we apply random search on a suffix to maximize a target logprob (e.g., of the token “Sure”), potentially with multiple restarts. In this way, we achieve nearly 100% attack success rate—according to GPT-4 as a judge—on Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4, and R2D2 from HarmBench that was adversarially trained against the GCG attack. We also show how to jailbreak all Claude models—that do not expose logprobs—via either a transfer or prefilling attack with a 100% success rate. In addition, we show how to use random search on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—which is the algorithm that brought us the first place in the SaTML’24 Trojan Detection Competition. The common theme behind these attacks is that adaptivity is crucial: different models are vulnerable to different prompting templates (e.g., R2D2 is very sensitive to in-context learning prompts), some models have unique vulnerabilities based on their APIs (e.g., prefilling for Claude), and in some settings, it is crucial to restrict the token search space based on prior knowledge (e.g., for trojan detection). For reproducibility purposes, we provide the code, logs, and jailbreak artifacts in the JailbreakBench format at https://github.com/tml-epfl/llm-adaptive-attacks.
@@ 1 Introduction
Llama-3-Instruct-8B Meta Full Prompt + random search + self-transfer None 100%
@@
The remarkable capabilities of Large Language Models (LLMs) carry the inherent risk of misuse, such as producing toxic content, spreading misinformation or supporting harmful activities. To mitigate these risks, safety alignment or refusal training is commonly employed—a fine-tuning phase where models are guided to generate responses judged safe by humans and to refuse responses to potentially harmful queries (Bai et al., 2022; Touvron et al., 2023). Although safety alignment is effective in general, several works have shown that it can be circumvented using adversarial prompts. These are inputs specifically designed to induce harmful responses from the model, a practice known as jailbreaking attacks (Mowshowitz, 2022; Zou et al., 2023; Chao et al., 2023).
@@
Jailbreaking attacks vary in their knowledge of the target LLM (ranging from white- to black-box approaches, or API-only access), complexity (involving manual prompting, standard optimization techniques, or auxiliary LLMs), and computational cost. Moreover, the nature of the jailbreaks they produce differs: some methods insert strings with little or no semantic meaning (Zou et al., 2023), while others rephrase user requests to maintain natural language (Mehrotra et al., 2023). The effectiveness of these attacks can significantly vary, achieving a high success rate on some target models but also drastically failing on others. Finally, some LLMs, such as the Llama-2-Chat family (Touvron et al., 2023), seem to maintain their robustness against these attacks. At the same time, new defensive mechanisms designed to counteract jailbreaks are emerging (Robey et al., 2023; Mazeika et al., 2024).
@@
In this work, we examine the safety of leading safety-aligned LLMs in terms of robustness to jailbreaks. We show that it is feasible to leverage the information available about each model, derived from training details or inference (e.g., logprobs), to construct simple adaptive attacks. Our main tool consists of a manually designed prompt template—which is used for all unsafe requests for a given model—enhanced by an adversarial suffix found with random search (RS) (Rastrigin, 1963) when the logprobs of the generated tokens are at least partially accessible. Our approach can be considered simple as it does not require gradient information (even for open-weight models) (Zou et al., 2023; Geisler et al., 2024), auxiliary LLMs to iteratively optimize the jailbreaks (Chao et al., 2023; Mehrotra et al., 2023; Zeng et al., 2024), or multi-turn conversations (Cheng et al., 2024; Russinovich et al., 2024). In this way, using the dataset of 50 harmful requests from AdvBench (Zou et al., 2023) curated by Chao et al. (2023), we obtain a close to 100% attack success rate on all leading safety-aligned LLMs, including Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4, Claude-3, and the adversarially trained R2D2. A summary of our main results in Table 1 suggests that our methods substantially outperform the existing attacks and achieve a 100% attacks success rate on many models for the first time. Additionally, we summarize all our evaluations in Table 22 in the appendix. We also show an illustrative example of a successful transfer attack on Claude 3 Sonnet in Figure 1. Finally, we show how to use RS on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—enabling us to secure the first place in the SaTML’24 Trojan Detection Competition (Rando & Tramèr, 2024).
@@
Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it is evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020); Croce et al. (2022b), distilling key observations from our experiments.
@@ 3.1 Setting
We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 9 for the prompt) in line with the criteria established by Chao et al. (2023), where a jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. We also include results using the rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6 for comparison. Additionally, we manually inspect all generations and flag cases with a significant number of false positives (which we only observed on Claude 2.1).
@@ 3.2 Methodology
Prior works define adaptive attacks as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). We follow this definition and describe the building blocks of our adaptive attacks, which we combine and potentially adapt depending on the target LLMs.
@@
The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We have optimized the set of rules one by one on the GPT-3.5 Turbo model to maximize the attack success rate and avoid the built-in safety guardrails. We provide the complete template in Table 2. As we will see in Section 4 (Table 5), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We have also designed another prompt template (referred to as the in-context prompt) that includes an example of unsafe behavior which the model is encouraged to imitate (see Table 8).
@@
We use a simple random search (RS) algorithm (Rastrigin, 1963) adapted for jailbreaking language models. The algorithm can be summarized as follows:
@@
We use adversarial suffixes initialized with $25$ tokens (which we found to be a nearly optimal number—see an example on Gemma-7B in Appendix B.1), we use up to 10 0001000010\,00010 000 iterations and up to 10 random restarts (although in most cases a single restart suffices). This strategy, including the preference for suffixes over prefixes and the focus on maximizing the probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We also tried other affirmative words, such as “exactly” and “certainly”, but we did not observe any improvement in the attack success rate. We opted for random search (a) due to its simplicity and efficiency, requiring only logprobs instead of gradients (thus also reducing the memory demands), and (b) motivated by its success in attacking vision models (Andriushchenko et al., 2020; Croce et al., 2022a).
@@
It is well-known that initialization plays a key role in optimization algorithms, including in RS-based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request as the initialization for RS on more challenging requests. We refer to this approach as self-transfer. Interestingly, this approach often leads to a transferable adversarial suffix (or, at least, a good starting for a subsequent run of random search), even though it is crafted for a single model and a single request. It significantly boosts the query efficiency and attack success rate on many LLMs.
@@
Some APIs like Claude allow users to directly prefill the LLM’s response with a specified beginning, making the aforementioned optimization procedure unnecessary. In that case, we explore prefilling the response with a string that corresponds to the target behavior (e.g., “Sure, here is how to make a bomb”). As a side note, the prefilling attack is also straightforward to implement for any open-weight LLM where the conversation template can be directly modified (Vega et al., 2023).
Reproducibility.
For reproducibility purposes, we provide all further details in our repository https://github.com/tml-epfl/llm-adaptive-attacks that contains the code, logs, and jailbreak artifacts in the JailbreakBench format.
@@ 4 Jailbreaking Leading Safety-Aligned LLMs
In this section, we detail the adaptive attacks we have developed for several families of leading safety-aligned LLMs. We provide a detailed descriptions of the main evaluations here and show the rest in Table 22 in the appendix where we also present results on Vicuna-13B, Mistral-7B, Phi-3-Mini, and Nemotron-4-340B.
4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
Here, we focus on open-weights Llama-2-Chat (7B, 13B, 70B parameters) (Touvron et al., 2023), Llama-3-Instruct (8B parameters) and Gemma-7B models (Google, 2023). These models have been significantly safety-trained, rendering them resilient to jailbreaks even in white-box scenarios (Zou et al., 2023).
@@ 4.1 Jailbreaking Llama-2, Llama-3, and Gemma Models
The key element to jailbreak the Llama models is self-transfer, where successful adversarial suffixes found by RS on simpler requests are used as initialization for RS on more complex requests. Notably, these adversarial strings tend to be to some extent transferable across different model sizes (e.g., from 7B to 13B models), but for the best result we repeat the self-transfer procedure for each model size separately. The same approach is also successful on Gemma-7B, although prompt + RS alone already demonstrates high attack success rate.
@@ Results.
Table 3 shows that we achieve 100% attack success rate on all these models. For Llama-2-Chat models, our standard adversarial prompt templates yield a 0% success rate, confirming the effectiveness of their safety alignment. When we apply Prompt + RS the attack success rate (ASR) increases to 48%. Ultimately, our composite attack strategy—which combines prompting, random search, and self-transfer—achieves a 100% attack success rate for all LLMs, surpassing all existing methods. For Llama-2-Chat-7B, the best reported success rate is 92% by PAP (Zeng et al., 2024) which is an LLM-assisted method. However, this method requires 10 restarts to achieve such accuracy, and its success rate drops to 46% with only one restart. In contrast, for this model, one restart is sufficient for our method to achieve a 100% ASR. Meanwhile, for the 13B and 70B models, Mazeika et al. (2024) reports ASR below 40%, while there is no prior evaluation available for Llama-3-Instruct and Gemma-7B since they are relatively recent models.
Table 3: Llama-2, Llama-3, and Gemma models. We report the attack success rate using the GPT-4 judge.
@@
Llama-3-Instruct-8B Prompt Ours 0%
Llama-3-Instruct-8B Prompt + random search Ours 100%
Llama-3-Instruct-8B Prompt + random search + self-transfer Ours 100%
@@
Convergence plots.
We show convergence curves in Figure 2, where we plot the average logprob of the token ‘Sure’ and average attack success rate for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) for random search (RS) with and without self-transfer. The plot confirms that starting from a good initialization via self-transfer is key for query efficiency and high success rate.
Figure 2: Convergence curves. We show the average logprob of the token ’Sure’ and attack success rate for three representative models (Llama-3-Instruct-8B, Llama-2-Chat-7B, and Gemma-7B) with and without self-transfer. Starting from a good initialization via self-transfer is key for query efficiency and high attack success rate.
@@ 4.2 Jailbreaking R2D2 Model
R2D2 uses adversarial training (Madry et al., 2018), a technique effective for obtaining vision models robust to $\ell_{p}$-bounded adversarial perturbations (Madry et al., 2018; Croce et al., 2021), to make LLMs more robust to jailbreak attacks.
@@
Similarly to Llama-2-Chat, the standard prompt template, alone or with RS, shows limited effectiveness. However, in contrast to Llama-2-Chat, self-transfer is ineffective here. Motivated by the fact that R2D2 was trained to refuse a specific prompt structure (i.e., request and adversarial suffix), we circumvent safety guardrails by using a different prompt structure. We use an in-context learning prompt (see Table 8 in the appendix), which we found the model to be particularly sensitive to. We use random search on top of the in-context prompt to maximize the probability of the initial token “Step” (instead of “Sure”) to be consistent with the new prompt template.
@@ Results.
The limitation of the API providing only the top-5 log-probabilities is not critical, as it is often straightforward to prompt a desired token, like “Sure”, to appear in the top-5. A more challenging issue is the non-deterministic output, since RS does not necessarily have a correct signal to refine the adversarial string. As illustrated in Figure 3, identical queries can yield varying log-probabilities, even with a fixed seed parameter and temperature zero in the API. The randomness makes random search less effective, although it still succeeds to a large extent.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 4.4 Jailbreaking Claude Models
Claude models are known for their high safety levels. In line with this, Anthropic does not provide access to logprobs for these models which prevents direct iterative attack like random search. Thus, we first test a transfer attack using an adversarial suffix optimized on GPT-4 with random search. We enhance the attack with multiple random restarts to leverage different generations with temperature one. Then we investigate an attack method that utilizes Anthropic’s prefilling feature, [^1] https://docs.anthropic.com/claude/docs/prefill-claudes-response which is not commonly available from other LLM API providers like OpenAI. To improve the attack success rate, we use the prefilling feature together with our prompt (Table 2) which we split into a system and user part (see Tables 17 and 18 for a detailed ablation).
@@
The prefilling feature makes jailbreaking straightforward on Claude models, even without any search (Table 6). For comparison, the previous best result on Claude 2.0 is 61% (Shah et al., 2023) while we get 100% using only up to 10 random restarts. The latest Claude 2.1 model, released in November 2023, is significantly more robust to both transfer and prefilling attacks. Nonetheless, we are able to get $100\%$ ASR with 100 restarts. We note that GPT-4 as a semantic judge sometimes has false positives, more often so on Claude 2.1. At the same time, the technical report of Claude-3 (Anthropic, 2024) mentions fewer refusals in its release announcement, which, in our findings, correlated with vulnerability to jailbreaking. We provide more complete experimental results, including the number of restarts in each case, in Tables 17 and 18 in the appendix.